b) A news website reported (Feb. 17, 2020) that ridership of Karachi Metro Bus declined after a fare increase: “There were nearly four million fewer riders in December 2019, the first full month after the price of a token increased Rs. 25 to Rs.150, than in the previous December, a 4.3 percent decline.”
i. Use these data to estimate the price elasticity of demand for Karachi Metro Bus. (2 Marks)
ii. According to your estimate, what happens to the Karachi Metro Bus’s revenue when the fare rises? (0.5 Marks)
iii. Why might your estimate of the elasticity be unreliable? (0.5 Marks
1
Expert's answer
2020-11-30T16:38:16-0500

i) Price elasticity of demand

= "\dfrac {\% \space change \space in \space Quantity}{\% \space change \space in \space price}"


"\% \space change \space in \space price = \dfrac {Rs. 150-Rs.25}{Rs. 25} × 100\%"


"= \dfrac {Rs. 125}{Rs. 25}×100\%"


"= 500\%"



"Thus, \space \eta_d = \dfrac {4.3\%}{500\%}"


"= 0.0046"



"\therefore \bold {\space PED = 0.0046}"



ii) When fare rises, the revenue for Karachi Metro Bus will rise as well. The rise in revenue is a result of the inelastic demand for the buses. Quantity demanded responds by less than proportionate to a change in price, causing revenue increase when price increases.


c) Because the fall in quantity demanded was associated with the increase in price, ignoring other factors affecting demand. It is possible that the fall in quantity demanded and an increase in price can be coincidental.